On July 1, 2016, Alpha Company exchanged an old computer (Equipment) with a historical cost of $1,000 that had accumulated depreciation of $600 after all June adjusting entries had been processed. The exchange was for a new computer having a fair value of $500. The transaction has commercial substance. Using this information, how much should be recorded on July 1 for the following accounts:

Accumulated Depreciation, Equipment

Gain or (Loss) on Sale (Enter any loss amount with $ sign inside of brackets)

Equipment - New
Answer #1
Accumulated Depreciation, Equipment 600
Gain or (Loss) on Sale 100 =500-(1000-600)
Equipment - New 500
When the transaction has commercial substance, the asset exchanged is recorded at fair value
Answer #2

Answer

--[1]Accumulated Depreciation would be debited by the amount of $ 600.

--Book Value of asset given up = 1000 - 600 = $ 400
--fair value of asset received = $ 500

--Hence, there is a Gain on $ 100 [500 - 400] on this exchange transaction.

--[2]Gain on Sale will be CREDITED by $ 100

--[3]Equipment (new) will be debited by fair value of $ 500, and Equipment (old) will be credited by the cost which was $ 1000